OpenStudy (anonymous):

You can model the population of a certain city between the years 1955 and 2000 by the radical function P(x)=55000 sqrt x-1945 . Using this model, in what year was the population of that city 275,000?

OpenStudy (anonymous):

\[275000=55000\sqrt{x-1945}\]

OpenStudy (anonymous):

Wait the 1945 is under the radical too right

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Ok, divide both sides by 55000 \[5=\sqrt{x-1945}\]

OpenStudy (anonymous):

square both sides 5^2=x-1945

OpenStudy (anonymous):

add 1945 1970=x

OpenStudy (anonymous):

Wut

OpenStudy (anonymous):

1970?

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

thx